# Post-glitch Recovery and the Neutron Star Structure: The Vela Pulsar

### v1

@@ III The Vortex Creep and Bending Models
Hypothesis III – An oscillatory model as described in Gügercinoğlu et al. (2023) consisting of .
The evidence for these hypotheses was also evaluated using Bayesian analysis to select the most suitable model and calculate its parameters. The bending of vortices remains a relatively unexplored domain, yet it can significantly impact simulations of neutron star glitches. Incorporating oscillation parameters associated with vortex bending, as reported in the next section, may provide a more comprehensive understanding of glitch behavior and boost the hydrodynamic simulation of the superfluid vortices (Howitt et al., 2020; Anantharaman and Bhattacharya, 2025).
IV Results and Discussions
@@
Our pulsar monitoring program with the ORT and the uGMRT started in 2015. The previous results were reported in Basu et al. (2020) and Grover et al. (2024a). In this work, we present analysis of a new Vela glitch at MJD 60429.9. Further, we demonstrate the spin-down rate evolution for the Vela pulsar since the beginning of the monitoring program. Our primary focus is to provide the theoretical interpretation of glitches using the post-glitch recovery phase and to obtain measurements for the re-coupling timescales and fractional moment of inertia for various layers that participate in a glitch. We have estimated the theoretical time to the next glitch through the vortex creep model and used Bayesian statistics to predict the time for the next glitch. All the predictions match well with the observed inter-glitch times.
IV.1 Glitches
@@
Here, we summarize the last three glitches in the Vela pulsar: MJD 57734.4, 58517, and 59417.6, detected in our monitoring program (Grover et al., 2024a; Basu et al., 2020). The recoveries for each of these glitches are studied within the framework of the vortex creep model to investigate the structure of the Vela pulsar. Furthermore, we report the detection and analysis of the recent large glitch in the Vela pulsar in 2024. The characteristics and parameters of glitches are listed in Table 1.
@@
| G1 | 57734.4(2) | 11.186433252(6) | -1.556383(8) | 1433.2(9) | 5.595(9) | Basu et al. (2020) |
| G2 | 58517(7) | 11.185397242(3) | –1.555659(4) | 2471(6) | 6(2) | Grover et al. (2024a) |
| G3 | 59417.6(1) | 11.184208478(5) | –1.5550(8) | 1235(5) | 8.0(7) | Grover et al. (2024a) |
@@
We present the parameters of the Vela 2024 glitch, which occurred around MJD 60429.9. This glitch has been reported by several groups (Campbell-Wilson et al., 2025; Zubieta et al., 2025b). Here, we present updates in our preliminary results reported in Grover et al. (2024b). The glitch is estimated to occur at MJD 60429.9(1), with the fractional increase in the spin frequency, and its time derivative is calculated as and , respectively. The results of glitch epoch and spin parameters agree with the reported estimates of other monitoring programs. A more precise measurement of glitch epoch as 60429.86962(4) is also presented in Palfreyman (2024) with higher cadence observations. The timing residuals, the evolution of the spin frequency and the spin-down rate are given in Fig. 2 and are being presented for the first time for this new glitch. The post-glitch recovery for this glitch is also investigated within the framework of the vortex creep model in the next subsection.
IV.2 Post-Glitch Recoveries
In this section, we present the post-glitch recovery behavior of the Vela pulsar. The spin-down rate has been analyzed from MJD 57651 to 60700. During this period, the Vela pulsar has displayed four glitches. The post-glitch recovery of the Vela pulsar is regular, i.e., it follows similar behavior. The recovery for this pulsar usually consists of exponential and linear relaxations well described by Model 3. We have investigated all post-glitch recoveries for this pulsar within the framework of the vortex creep model for the glitches listed in Table 1, and the resulting fit parameters obtained using the vortex creep model are listed in Table 2. Now, we concisely describe the post-glitch relaxation(s) for individual cases.
@@ IV.2 Post-Glitch Recoveries
G1, or glitch at MJD 57734.4(2) (Palfreyman et al., 2018; Basu et al., 2020), was detected with a fractional change in the rotational frequency of 1433(1), and a fractional change in the spin-down rate as 5.59(1). We present the post-glitch relaxation modeling for this glitch in Fig. 3. For this recovery, the Bayesian model comparison prefers the presence of two exponentials (Model 3b) over all other models. The estimated parameters are: the fractional moment of inertia for the second and third exponential components are and respectively, with the decay timescales of and days, respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of 61 days. The offset time is measured as days, close to the observed inter-glitch time of 783 days. The observed time is consistent with the vortex creep model’s predicted range and closely matches the Bayesian prediction, which estimates a 68% credible interval of 787 to 1083 days. The fit parameters are listed in Table 2.
@@
The bottom part of Fig. 3 shows the residuals obtained after subtracting the vortex creep model predictions from the observed data and is in units of Hz s-1. These vortex residuals, spanning from 200 days onward, were utilized for vortex bending studies, excluding the initial 200 days due to the significant impact of exponential recovery. The generalized Lomb-Scargle periodogram analysis, Fig. 4, revealed a prominent peak corresponding to a period of 314.1(2) days, with a false alarm probability (FAP) of (Z-score: 4.9). A secondary peak at approximately 166 days had a FAP of 0.194, and other subsequent peaks had FAP values of 1, indicating low significance. Fixing the period at 314.1(2) days, we investigated the vortex bending hypothesis and found that the damping model, comprising a sinusoidal term and a damping factor (Hypothesis II), was strongly favored over Hypotheses I and III. Additionally, we examined whether the observed peak resulted from shocks caused by the current glitch or the previous one, concluding that the oscillations decisively originated from the current glitch. The comparison of vortex residuals and vortex bending model is shown in Fig. 5 and the estimated parameters for the damped model (Hypothesis II) are: Amplitude = , Phase = , and days. Note that vortex residuals are in scale of Hz s-1, so the effective amplitude is .
The vortex residuals fit with the vortex bending model well for around 200-400 days post-glitch duration but deviate thereafter. Including a second peak improves the fit, but the generalized Lomb-Scargle periodogram indicates a FAP of 19.4% for this peak using the Baluev method (with a minimum FAP of 8.9% obtained through approximations based on effective independent frequencies). Given the relatively high FAP, we exclude the second component from our analysis. The other peaks with a high FAP are likely to originate from preceding glitches, and their effect on the present glitch recovery is insignificant. The post-glitch recovery modeling for this pulsar is also presented by Gügercinoğlu et al. (2022) using the Fermi-LAT observations. However, they only detected one exponential component. We are reporting 2 exponential recoveries made possible due to much higher cadence observations allowing us to investigate the vortex residuals in the framework of the vortex bending model for the first time.
IV.2.2 Glitch No. 2 – MJD 58517(7)
The glitch at MJD 58517(7), G2 in the Vela pulsar, (Grover et al., 2024a; Lower et al., 2020; Gügercinoğlu et al., 2022), was observed with a fractional change in the rotational frequency of 2471(6), and a fractional change in the spin-down rate as 6(2). Our analysis focuses on modeling the post-glitch relaxation behavior, given in Fig. 6. Observations weren’t conducted during the exponential relaxation phase of this glitch, and we detected only non-linear creep recovery. To proceed with our post-glitch analysis, we supplemented our dataset with publicly available ATNF pulsar data to fill temporal gaps. Additionally, we utilized estimates derived from analysis using data closer to the glitch from Lower et al. (2020), specifically the glitch epoch (58515.6) and the fractional change in rotational frequency and it’s derivative (2501 and 8.7, respectively), for our recovery studies.
The Bayesian model comparison reveals that the three exponential model (Model 3c) best describes this recovery. The measured values for the recovery parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The offset time is estimated as days. The observed inter-glitch time of 902 days is consistent with both the predicted range of the vortex creep model and Bayesian statistics. The median value predicted by Bayesian statistics is 893 days, very close to the observed time. These parameters are also summarized in Table 2.
The first exponential recovery timescale is approximately 7 days, roughly 3 or more times larger than the corresponding component in the next two glitches. Given this disparity, we investigated the possibility of an additional exponential with a shorter timescale (0.1–2 days) but found no converging results, suggesting that the 7-day timescale is a distinct characteristic of the initial recovery phase. The third exponential component’s precision is biased due to poor constraint, affecting the measurements of vortex residuals. The third exponential recovery’s impact on vortex bending is its timescale of approximately 600 days; hence, we did not study vortex residuals for this glitch using the vortex bending model. Our post-glitch recovery modeling results differ from those of Gügercinoğlu et al. (2022), who reported a single exponential component using Fermi-LAT observations. Our analysis benefits from a higher cadence and reveals two well-constrained distinct exponential recoveries.
IV.2.3 Glitch No. 3 – MJD 59417.6(1)
G3, or the glitch at MJD 59417.6(1) (Grover et al., 2024a; Zubieta et al., 2023) was observed to have a fractional spin up in frequency and a fractional spin up in frequency-time derivative . We present the post-glitch relaxation behavior for this glitch in Fig. 7. In addition to the ORT, we leverage observed data from the uGMRT as well, with two motivations: (i) to confirm that the residuals stem from the neutron star’s interior, rather than the ionized interstellar medium by utilizing DM offsets in the uGMRT analysis and (ii) to fill the temporal gap between 850 to 1000 days subsequent to the glitch date.
High-cadence ORT observations of this glitch enable us to model three exponential components present in the recovery; Model 3c emerged as the top choice in Bayesian model comparisons, outperforming all other models. The calculated values for the parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days, respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The offset time is estimated as days, which deviates significantly from the observed inter-glitch time primary due to the large value of (equivalent to ). However, the vortex creep model predicted time for the next glitch lies between to , which aligns well with the observed time. The median Bayesian prediction of 989 days is near the observed interval of 1012 days, falling well within the 68% credible interval. These measured parameters are listed in Table 2.
The relaxation behavior of this glitch has never been investigated before. The lower panel of Fig. 7 shows the vortex residuals, obtained after subtracting the vortex creep model predictions from the observed data, in scale of Hz s-1. We analyzed residuals between 200 and 700 days for vortex bending studies, excluding the initial 200 days due to dominant exponential recovery effects and data beyond 700 days, which warrants deep investigation.
@@
The Lomb-Scargle periodogram analysis of vortex residuals, depicted in Fig. 8, exhibits multiple peaks. However, only the first two peaks, corresponding to periods of approximately 344(6) and 153(3) days, demonstrate statistical significance. The first peak yields a FAP of 4.6 10-13 and a Z-score of 7.1, while the latter peak has a FAP of 2.46 10-5 and a Z-score of 4.1. The third peak, with a period of around 81 days, has a FAP of 0.88 and is not significant, nor other peaks all of which have a FAP of 1. These two peaks are used to investigate the vortex residuals within the framework of the vortex bending hypothesis, presented in Fig. 9. We observed that Hypothesis II was strongly favored over Hypotheses I and III. Additionally, we examined whether the observed peak resulted from shocks caused by the current glitch or the previous one, concluding that the oscillations decisively originated from the current glitch. The estimated parameters for the damped model (Hypothesis II) are: (i) Amplitudes, A1 = , and A2 =, (ii) Phase, and , (iii) and days. Note that vortex residuals are in scale of Hz s-1, so the effective amplitudes are , and , respectively.
IV.2.4 Glitch No. 4 – MJD 60429.9(1)
We present the post-glitch relaxation, presented for the first time, for the recent glitch at MJD 60429.9(2), G4 (Grover et al., 2024b; Campbell-Wilson et al., 2025; Zubieta et al., 2025b). The glitch plot is given in Fig. 2, and the comparison of the observed data with the vortex creep model fitting is presented in Fig. 10.
@@
The Bayesian model comparisons selected the Model 3c as the best model over all other models. The calculated values for the parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days, respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The time for the next glitch is predicted as days. These parameters are listed in Table 2. The parameters for the linear creep regions (or exponential recovery) are well-constrained. However, those dependent on the non-linear creep region will likely be refined in the future as more data becomes available. The Bayesian analysis forecasts a median time for the next glitch at MJD 61377.7, with a 68% credible interval spanning from MJD 61249 to 61506.
IV.3 Correlations and Braking Index

### v2

@@ 2 The Vortex Creep and Bending Models
Hypothesis III – An oscillatory model as described in Erbil_2023 consisting of .
where is the amplitude, is the angular frequency, is the phase, and is the decay time for the oscillation. The next section outlines the observations, data reduction, and timing used to derive the spin frequency and spin-down rate. It also describes the Bayesian analysis used to calculate the evidence for the competing hypotheses, thereby identifying the most plausible model and estimating its corresponding parameters.
3 Observations and Analysis
3.1 Sample and Observational Facilities
We use two large radio telescopes to monitor pulsars: the upgraded Giant Metrewave Radio Telescope (uGMRT)(ugmrt2017) and the Ooty Radio Telescope (ORT) (swarup1971ort). Observations of PSR B083345 were part of a regular glitch monitoring program, where we regularly monitor 24 pulsars (i) to detect glitches, (ii) model timing noise, and (iii) model post-glitch recovery. This work presents a comprehensive analysis of the Vela pulsar’s rotational evolution, spanning from September 2016 to January 2025, as part of our ongoing monitoring campaign (Grover2024; Basu2020). Our study primarily uses observations from the ORT, supplemented by uGMRT and archival ATNF data to fill temporal gaps and facilitate specialised investigations.
The ORT (swarup1971ort) is a 530m 30m parabolic cylindrical reflector. It is located on a hill with the same slope as the geographic latitude (11°), allowing it to track sources for about 8 hours and covers a declination range of -60° to +60°. Pulsar observations at the ORT are carried out at the central observation frequency of 326.5 MHz using the Pulsar Ooty Radio Telescope’s New Digital Efficient Receiver, PONDER (Naidu_2015), which produces real-time coherent dedispersed time series data. The typical cadence of the observations at the ORT is 1–14 days.
The GMRT is an interferometer with 30 parabolic antennas, with a diameter of 45 meters each (Swarup1991GMRT; Ananthakrishnan1995GMRT). The uGMRT (ugmrt2017) covers a declination range of -53°to +90°and supports observations from 150 MHz to 1460 MHz in four bands. The Vela pulsar is observed in uGMRT Band 3 (250–500 MHz) and Band 4 (550–950 MHz). The central square antennas, together with a subset of the nearest arm antennas, were used in a phased array to synthesise a large single dish for observations. The outer arm antennas were not included in the phased array due to significant drift in the phase of the voltage outputs, caused by ionospheric variations, in these antennas. A digital filterbank, implemented in the GMRT correlator, was used to record either 1024 or 2048 channels across the bandpass to obtain frequency-resolved data with a typical sampling time of 655 s. The cadence of the observations is 15-30 days. We primarily use frequency-resolved uGMRT data to examine the impact of dispersion measure (DM) offsets on the spin-down rate. Additionally, uGMRT data filled a temporal gap around epoch 60300, providing crucial estimates not covered by ORT due to maintenance downtime.
The Commonwealth Scientific and Industrial Research Organisation (CSIRO) Data Access Portal is used to access the Parkes pulsar archival observations222https://data.csiro.au/domain/atnf (Hobbs2011) of the Vela pulsar. These data were mainly used to fill temporal gaps in glitch number 2 (MJD 58515), caused by significant observational gaps at ORT. We directly downloaded and used the reduced files (or PSRFITS: PSRCHIVE_and_PSRFITS2004) and performed subsequent analysis, as described in the next subsection.
3.2 Data Reduction and Timing Analysis
We used the DSPSR software (DSPSR) and the PINTA pipeline (pinta) to reduce the raw data from the Vela pulsar observed at the ORT and the uGMRT, respectively, into PSRFITS files, using the ephemeris created in our initial observations. The initial analysis, including the generation of a noise-free template and Time of Arrivals (ToAs), was performed using PSRCHIVE (PSRCHIVE_and_PSRFITS2004; PSRCHIVE2012). The ToAs were obtained through the cross-correlation of the template profile with all other observed profiles using the ‘pat’ command, using the Fourier Phase Gradient (PGS) method (Taylor1992). The precision timing analysis of the pulsar was done with the help of the pulsar timing software TEMPO2 (Hobbs2006Tempo2; Edwards2006Tempo2). The DM offsets were estimated using the DMcalc (Krishnakumar2021) script from the frequency-resolved data from uGMRT. The rotational evolution of a pulsar including a glitch can be expressed as a Taylor series expansion (Erbil_2022),
where is the spin frequency, and and are the first and second derivatives of the frequency, respectively, at reference epoch . and denote the changes in spin frequency and its first derivative at the glitch epoch , respectively. The parameters and correspond to the amplitude and characteristic decay timescale of the exponential relaxation component following the glitch.
The spin evolution of the Vela pulsar from September 2016 to January 2025, obtained after precision timing, is given in Fig. 1. Each measurement is made with the local timing of several epochs, typically within a 50-200 day data window, contingent on data availability. Exceptions include the first and last few points adjacent to glitches. The resulting measurements are reflected in the precision of error bars, which are smaller than the symbol used in the figure. By choosing multi-epoch windows, we achieve reliable rotational estimates with minimised error bars, providing a robust foundation for studying vortex residuals. This approach effectively averages out temporal effects caused by physical processes, such as ionised interstellar medium scattering-induced broadening, thereby validating the accuracy of measurements aligned with the star’s rotation and internal dynamics. These recovery measurements are investigated within the framework of the vortex creep and bending models as described in Section 2 and the results are discussed in Section 4.
3.3 Bayesian Analysis
For each recovery, we used Bayesian statistics for model comparison and parameter estimation. Bayesian statistics use Bayes’ theorem to enhance the probability of a hypothesis based on new evidence or data. We used DYNESTY (a dynamic nested sampling package for estimating Bayesian posteriors and evidences; (DYNESTY2020)) to compute the Bayesian evidence and perform model selection, and EMCEE (EMCEE2013) to obtain posterior samples for parameter estimation of the preferred model. For model comparison, the standard Jeffreys1961 scale is used, which suggests the preferred model based on the strength of evidence (Trotta2008). Bayesian inference is also used to predict the epoch of the next glitch using data from previous glitches.
The priors are selected based on the initial knowledge of each parameter, as informed by the literature (Erbil_2022; Flanagan1990; Alpar1989). This knowledge provides reasonable initial ranges that can be further refined or broadened based on the results. For instance, parameters such as the fractional moment of inertia ( or ) is generally positive and below unity. The recoupling timescale in exponential regions () depends on the number of exponential components (three-, two-, or one-component models). Typically, for the Vela pulsar, it has been observed that the first component lies within 0.1–10 days, with the second and third components approximately 10 and 100 times longer, respectively.
For the linear relaxation region, the prior for recoupling timescale () is around 10-300 days, and the offset time () is 300 to 1200, with the initial guess often chosen close to the observed time to the next glitch. For the Vortex bending studies, the prior on amplitude, phase, and decay constant were set as 0–1, 0–2, and 0–2000 days, respectively. All the priors were considered as uniform priors and listed in Table 1. While these priors adequately describe many recoveries, particularly those observed in the Vela pulsar, their applicability may vary depending on the nature of the recovery. As each recovery exhibits unique characteristics, a flexible and adaptive approach to defining priors is essential. We model the data assuming independent Gaussian measurement uncertainties. The corresponding log-likelihood function is given by
where denotes the observed data, is the model prediction for parameters , and represents the corresponding measurement uncertainty.
| Parameters | Values | Parameters | Values |
| 0–1 | 0–1 | ||
| 0–1 | 10–300 days | ||
| 0–1 | 300–1200 days | ||
| 0.1–10 days | 0–1 | ||
| 5–100 days | 0–2 | ||
| 10–400 days | 0–2000 days |
The post-glitch recoveries are studied within the framework of the vortex creep model and the vortex bending model (Erbil_2023), discussed in the previous section. A quick picture of the analysis process is as follows: We first model the frequency derivative measurements from the data using the vortex creep model. Then, we estimate the residuals obtained by subtracting the predicted values from the vortex creep model from the measured values. These residuals were investigated within the framework of the vortex bending model. The generalised Lomb-Scargle periodogram technique (Lomb1976; Scargle1982; VanderPlas2018) was employed to detect periodicities in the residuals. We used the Astropy (Astropy2013) implementation of the generalised Lomb-Scargle periodogram. The Peaks with false alarm probabilities (FAP) below 0.1% or Z-score above 3 were considered significant for vortex bending studies. FAP estimation was performed using the Baluev2008 method. Alternative methods, including approximations based on effective independent frequencies and bootstrap resampling, yielded similar results. The significance of each detected peak is quantified using the Lomb–Scargle false-alarm probability (FAP). For interpretive purposes, we also report the equivalent z-score values (Cowan2011; Ganguly_Desai_2017). The results of our analysis are presented in the next section.
4 Results and Discussions
@@
Our pulsar monitoring program with the ORT and the uGMRT started in 2015. The previous results were reported in Basu2020 and Grover2024. In this work, we present an analysis of a new Vela glitch at MJD 60429.9. Further, we demonstrate the spin-down rate evolution for the Vela pulsar since the beginning of the monitoring program. Our primary focus is to provide the theoretical interpretation of glitches using the post-glitch recovery phase and to obtain measurements for the re-coupling timescales and fractional moment of inertia for various layers that participate in a glitch. We have estimated the theoretical time to the next glitch through the vortex creep model and used Bayesian inference to predict the time for the next glitch. All the predictions match well with the observed inter-glitch times.
4.1 Glitches
@@
Here, we summarise the last three glitches in the Vela pulsar: MJD 57734.4, 58517, and 59417.6, detected in our monitoring program (Grover2024; Basu2020). The recoveries for each of these glitches are studied within the framework of the vortex creep model to investigate the structure of the Vela pulsar. Furthermore, we report the detection and analysis of the recent large glitch in the Vela pulsar in 2024. The characteristics and parameters of glitches are listed in Table 2.
@@
| G1 | 57734.4(2) | 11.186433252(6) | -1.556383(8) | 1433.2(9) | 5.595(9) | Basu2020 |
| G2 | 58517(7) | 11.185397242(3) | –1.555659(4) | 2471(6) | 6(2) | Grover2024 |
| G3 | 59417.6(1) | 11.184208478(5) | –1.5550(8) | 1235(5) | 8.0(7) | Grover2024 |
@@
We present the parameters of the Vela 2024 glitch, which occurred around MJD 60429.9. This glitch has been reported by several groups (Zubieta2024ATel16608_Vela; Campbell-Wilson2024ATel16610_Vela; Grover2024ATel16611_Vela; Palfreyman2024ATel16615_Vela; Wang2024ATel16619_Vela; Zubieta2025_Vela). Here, we present updates in our preliminary results reported in Grover2024ATel16611_Vela. The glitch is estimated to occur at MJD 60429.9(1), with the fractional increase in the spin frequency, and its time derivative is calculated as and , respectively. The results of the glitch epoch and spin parameters agree with the reported estimates of other monitoring programs. A more precise measurement of glitch epoch as 60429.86962(4) is also presented in Palfreyman2024ATel16615_Vela with higher cadence observations. The timing residuals, the evolution of the spin frequency, and the spin-down rate are given in Fig. 2. The post-glitch recovery for this glitch is also investigated within the framework of the vortex creep model in the next subsection.
4.2 Post-Glitch Recoveries
In this section, we present the post-glitch recovery behaviour of the Vela pulsar. The spin-down rate has been analysed from MJD 57651 to 60700. During this period, the Vela pulsar has displayed four glitches. The post-glitch recovery of the Vela pulsar is regular, i.e., it follows similar behaviour. The recovery for this pulsar usually consists of exponential and linear relaxations well described by Model 3 of Section 2. We have investigated all post-glitch recoveries for this pulsar within the framework of the vortex creep model for the glitches listed in Table 2, and the resulting fit parameters obtained using the vortex creep model are listed in Table 3. Now, we concisely describe the post-glitch relaxation(s) for individual cases.
@@ 4.2 Post-Glitch Recoveries
G1, or glitch at MJD 57734.4(2) (Sarkissian2017; Palfreyman2018_Vela; Ashton2019; Basu2020; Erbil_2020; Montoli2020; Dunn2025), was detected with a fractional change in the rotational frequency of 1433(1), and a fractional change in the spin-down rate as 5.59(1). We present the post-glitch relaxation modelling for this glitch in Fig. 3. For this recovery, the Bayesian model comparison prefers the presence of three-exponential + linear recovery (Model 3c) over all other models. The estimated median values for parameters are: the fractional moment of inertia for the first, second, and third exponential components are , and respectively, with the decay timescales of , and days, respectively. The median fractional moment of inertia for the non-linear creep region is with a recoupling timescale of 62 days. The offset time is measured as days, close to the observed inter-glitch time of 783 days. The observed time is consistent with the vortex creep model’s predicted range and closely matches the Bayesian prediction, which estimates a 68% credible interval of 787 to 1083 days. The fit parameters, along with their 95% credible intervals, are listed in Table 3.
@@
The bottom part of Fig. 3 shows the residuals obtained after subtracting the vortex creep model predictions from the observed data and is in units of Hz s-1. These vortex residuals, spanning from 200 days onward, were used for vortex bending studies, excluding the initial 200 days due to the significant impact of exponential recoveries. The generalised Lomb–Scargle periodogram (Fig. 4) reveals a prominent peak corresponding to a period of 315(1) days using the least-squares interpretation of the periodogram (VanderPlas2018). This signal exhibits strong significance, with a false alarm probability (FAP) of (Z-score: 5.0). A secondary peak at approximately 165 days has a FAP of 0.14, while all remaining peaks yield FAP values of 1, indicating low significance.
@@
Fixing the period at 315(1) days, we investigated the vortex bending hypothesis and found that the damping model, comprising a sinusoidal term and a damping factor (Hypothesis II), was strongly favoured over Hypotheses I and III. Additionally, we examined whether the observed peak resulted from shocks caused by the current glitch or the previous one, concluding that the oscillations decisively originated from the current glitch. The comparison of vortex residuals and vortex bending model is shown in Fig. 5 and the estimated parameters for the damped model (Hypothesis II) are: Amplitude = , Phase = , and days. Note that vortex residuals are in scale of Hz s-1, so the effective amplitude is .
The vortex residuals fit with the vortex bending model well for around 200-400 days post-glitch duration, but deviate thereafter. Including a second peak improves the fit, but the generalised Lomb-Scargle periodogram indicates a FAP of 14% for this peak using the Baluev method (with a minimum FAP of 5.3% obtained through approximations based on effective independent frequencies). Given that the FAP of the second peak is greater than 0.1%, we exclude the second peak from our analysis. The other peaks with a high FAP are likely to originate from preceding glitches, and their effect on the present glitch recovery is insignificant. The post-glitch recovery modeling for this pulsar is also presented by Erbil_2022 using the Fermi-LAT observations. However, they only detected one exponential component. We are reporting three exponential recoveries made possible by much higher cadence observations, which allow us to investigate the vortex residuals within the framework of the vortex bending model.
4.2.2 Glitch No. 2 – MJD 58517(7)
The glitch at MJD 58517(7), G2 in the Vela pulsar, (Sarkissian2019ATel12466; Kerr2019ATel12481; LopezArmengol2019ATel12482; lower2020_utmost2; Erbil_2022; Grover2024; Dunn2025), was observed with a fractional change in the rotational frequency of 2471(6), and a fractional change in the spin-down rate as 6(2). Our analysis focuses on modelling the post-glitch relaxation behaviour, given in Fig. 6. Our ORT and uGMRT observations weren’t conducted during the initial exponential relaxation phase of this glitch, and we detected only non-linear creep recovery. To proceed with our post-glitch analysis, we supplemented our dataset with publicly available ATNF pulsar data to fill temporal gaps. Additionally, we used estimates derived from analysis using data closer to the glitch from lower2020_utmost2, specifically the glitch epoch (58515.6) and the fractional change in rotational frequency and its derivative (2501 and 8.7, respectively), for our recovery studies.
The Bayesian model comparison reveals that the model with two exponential terms and a linear recovery term (Model 3b) provides the best description of this recovery. The measured median values for the recovery parameters are as follows: the fractional moment of inertia for the second and third exponential components are and respectively, with the decay timescales of and days respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The offset time is estimated as days. The observed inter-glitch time of 902 days is consistent with both the predicted range of the vortex creep model and Bayesian statistics. The median value predicted by Bayesian statistics is 893 days, very close to the observed time. These parameters are also summarised in Table 3. Unlike other glitches that were observed with very high cadence using the ORT, this glitch was not monitored with such precision, and therefore, the vortex-bending oscillations are not well observed and studied in this work. Our post-glitch recovery modelling results differ from those of Erbil_2022, who reported a single exponential component using Fermi-LAT observations. Our analysis benefits from a higher cadence and reveals two well-constrained, distinct exponential recoveries.
4.2.3 Glitch No. 3 – MJD 59417.6(1)
G3, or the glitch at MJD 59417.6(1) (Sosa-Fiscella2021ATel14806; Dunn2021ATel14807; Olney2021ATel14808; Singha2021ATel14812; Zubieta_2023; Grover2024; Dunn2025) was observed to have a fractional spin up in frequency and a fractional spin up in frequency-time derivative . We present the post-glitch relaxation behaviour for this glitch in Fig. 7. In addition to the ORT, we leverage observed data from the uGMRT as well, with two motivations: (i) to confirm that the residuals stem from the neutron star’s interior, rather than the ionised interstellar medium by using DM offsets in the uGMRT analysis and (ii) to fill the temporal gap between 850 to 1000 days subsequent to the glitch date.
@@
High-cadence ORT observations of this glitch enable us to model three exponential components present in the recovery; Model 3c emerged as the top choice in Bayesian model comparisons, outperforming all other models. The calculated median values for the parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days, respectively. The fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The offset time is estimated as days, which deviates significantly from the observed inter-glitch time primarily due to the large value of (equivalent to ). However, the vortex creep model predicted time for the next glitch lies between to , which aligns well with the observed time. The median Bayesian prediction of 989 days is near the observed interval of 1012 days, falling well within the 68% credible interval. These measured parameters are listed in Table 3.
The relaxation behaviour of this glitch has never been investigated before. The lower panel of Fig. 7 shows the vortex residuals, obtained after subtracting the vortex creep model predictions from the observed data, in scale of Hz s-1. We analysed residuals between 200 and 700 days for vortex bending studies, excluding the initial 200 days due to dominant exponential recovery effects, and data beyond 700 days requires further investigation in existing theory.
The Lomb–Scargle periodogram analysis of the vortex residuals (Fig. 8) reveals several peaks, but only the first two are statistically significant. The strongest peak yields a FAP of 4.6 10-13 and a Z-score of 7.2, while the latter peak has a FAP of 2.46 10-5 and a Z-score of 4.1. The third peak, with a period of around 81 days, has a FAP of 0.88 and is therefore not significant; all remaining peaks have FAP values consistent with unity. The corresponding periods of the two significant peaks are approximately 344.0(7) and 153.5(7) days, estimated using the least-squares interpretation of the periodogram (VanderPlas2018).
These two peaks are used to investigate the vortex residuals within the framework of the vortex bending hypothesis, presented in Fig. 9. We observed that Hypothesis II was strongly favoured over Hypotheses I and III. Additionally, we examined whether the observed peak resulted from shocks caused by the current glitch or the previous one, concluding that the oscillations decisively originated from the current glitch. The estimated parameters for the damped model (Hypothesis II) are: (i) Amplitudes, A1 = , and A2 =, (ii) Phase, and , (iii) and days. Note that vortex residuals are in scale of Hz s-1, so the effective amplitudes are , and , respectively.
4.2.4 Glitch No. 4 – MJD 60429.9(1)
We present the post-glitch relaxation for the recent glitch at MJD 60429.9(2), G4 (Zubieta2024ATel16608_Vela; Campbell-Wilson2024ATel16610_Vela; Grover2024ATel16611_Vela; Palfreyman2024ATel16615_Vela; Wang2024ATel16619_Vela; Zubieta2025_Vela). The glitch is illustrated in Fig. 2, and the comparison of the observed data with the vortex creep model fitting is presented in Fig. 10.
The Bayesian model comparisons selected Model 3c as the best model over all other models. The calculated median values for the parameters are as follows: the fractional moment of inertia for the first, second and third exponential components are and and respectively, with the decay timescales of , and days, respectively. The median fractional moment of inertia for the non-linear creep region is with a recoupling timescale of days. The median value for the offset time is days. These parameters, along with their 95% credible intervals, are listed in Table 3. The parameters for the linear creep regions (or exponential recovery) are well-constrained. However, those dependent on the non-linear creep region will likely be refined in the future as more data becomes available. The Bayesian analysis forecasts a median time for the next glitch at MJD 61377.7, with a 68% credible interval spanning from MJD 61249 to 61506.
4.3 Correlations and Braking Index